Aus tiefer Not schrei ich zu dir (d–phrygisch)

M: Martin Luther 1524 • T: Martin Luther 1524 nach Psalm 130 • Choralsatz: 26. 2.2010 Norbert Südland

[figure: 1.; {Aus \atop dein}; {tie- \atop gnä-}; {fer \atop dig}; {Not \atop Oh-}; schrei ren; {ich \atop kehr}; {zu \atop zu}; {dir, \atop mir}; {Herr \atop und}; {Gott, \atop mei-}; {er- \atop ner}; {hör \atop Bitt}; {mein \atop sie}; {Ru- \atop öf-}]

[figure: 1.; fen;; 2.; fen!; }; Denn; so; du; willst; das; se-; hen; an,; was; Sünd; und]

Aus tiefer Not schrei ich zu dir (e–phrygisch)

M: Martin Luther 1524 • T: Martin Luther 1524 nach Psalm 130 • Choralsatz: 26. 2.2010 Norbert Südland

[figure: 1.; {Aus \atop dein}; {tie- \atop gnä-}; {fer \atop dig}; {Not \atop Oh-}; schrei ren; {ich \atop kehr}; {zu \atop zu}; {dir, \atop mir}; {Herr \atop und}; {Gott, \atop mei-}; {er- \atop ner}; {hör \atop Bitt}; {mein \atop sie}; {Ru- \atop öf-}]

[figure: 1.; fen;; 2.; fen!; }; Denn; so; du; willst; das; se-; hen; an,; was; Sünd; und]